OpenStudy (anonymous):

how to translate the graph y = |x| to obtain the graph of y = |x - 1| - 1?

OpenStudy (anonymous):

\[y=|x-1|\] moves \(y=|x|\) to the RIGHT one unit

OpenStudy (anonymous):

the \(-1\) out at the end moves it down one

OpenStudy (michele_laino):

Your equation can be rewritten as below: \[y+1=|x-1|\] so if we introduce the new coordinates X, Y defined as: X=x-1, and Y=y+1 your equation can be rewritten as below: \[Y=|X|\] which is in the desired form, so the equation of translation are:(1) \[X=x-1, Y=y+1\] now if in ewuations (1) we set X=0 and Y=0, namely the origin of the new system of coordinates, we get: \[x=1,y=-1\] namely you have to translate your coordinate system such that the new position of origin is located at point (1,-1)

OpenStudy (anonymous):

Thanks

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Which of the following describes how to translate the graph y = |x| to obtain the graph of y = |x + 7|?

OpenStudy (michele_laino):

please, you have to perform the subsequent translation: \[Y=y,X=x+7\] where X,Y are the new axes, in the new coordinate system your equation can be rewritten as below: \[Y=|X|\] so like before, your new origin is located at the point (-7,0)

OpenStudy (anonymous):

so would it be 7 units up or down or left or right?